OpenStudy (anonymous):

In which sentence is the fraction written correctly? A. The family took one-third of the toys to the donation center. B. We ate one-half of the pizza for dinner. C. Is 0.25 cup of melted butter the same as one-fourth cup of melted butter?

OpenStudy (anonymous):

i am not sure on this one

OpenStudy (anonymous):

i know c is wrong

iYuko (iyuko):

A and B look similar

OpenStudy (tinylittlehelper):

Okay, remember what we all said last time?

OpenStudy (anonymous):

so it is either a or b, i just can not decided

OpenStudy (anonymous):

together like friends

OpenStudy (anonymous):

but a and b do that

OpenStudy (tinylittlehelper):

Wait... with these, anything 1 and under has a singular noun following it. EX: I have 0.999999 piece of pie left. Want the slither?

OpenStudy (anonymous):

so a? i know c is wrong...

OpenStudy (anonymous):

I think its A. Not supposed to give out direct answers but I cant explain this sorry

OpenStudy (anonymous):

so it has got to be b then...

OpenStudy (tinylittlehelper):

Wait, yeah maybe...

OpenStudy (anonymous):

u r thinking it is a?

OpenStudy (tinylittlehelper):

Research...

OpenStudy (anonymous):

ok

OpenStudy (tinylittlehelper):

Date: 01/20/2004 at 13:15:46 From: Doctor Peterson Subject: Re: Writing fractions in word form Hi, Colleen. I usually reserve the hyphen for use in keeping the name of the denominator together. For example, 1/5 = one fifth 1/25 = one twenty-fifth 2/100,000 = two hundred-thousandths 200/1000 = two hundred thousandths But you will also see hyphens within numbers when the entire number is used in certain special settings: One third of a piece A one-third share A hundred thousand dollars A hundred-thousand-dollar income http://mathforum.org/library/drmath/view/64567.html

OpenStudy (anonymous):

does this mean it is a?

OpenStudy (tinylittlehelper):

One more to see... I think so, but this might be okay. Hyphens and Spelled-out Fractions Fractions should always be hyphenated when they are adjectives or adverbs, as in They got a one-third share and The money is three-quarters gone. Opinions differ on whether they should be hyphenated when they are nouns, as in They got one-third of the money. By standard rules of hyphenation, there is no reason to hyphenate them; they are merely noun compounds formed of adjective + noun. However, the hyphen is "heard" -- we do not pronounce the elements of such compounds as distinct words but slur them together--and omission of the hyphen could conceivably mislead: I used to save all of my change in a bucket, but I've spent three quarters of it. We suggest hyphenating fractions routinely. One can think of the hyphen as representing the division bar in a fraction in figures. The horse rounded the track five and three-quarters times. Adjectival compounds of a whole number and a fraction are not hyphenated throughout unless they are complicated by another word, as in The horse fell at the one-and-one-quarter-mile mark.

OpenStudy (anonymous):

so it has got to be a then, i am going with that @TinyLittleHelper , u r a life saver!

OpenStudy (tinylittlehelper):

Well, thanks; I kinda guessed with the research. Google comes in handy! :)